Metromarket
18 cans of soda for $9

Wal- town
10 cans of soda for $7.50
Shopstop
14 cans of soda for $7.35
Which store sell the best deal one can of soda? Explain

Answer by checkley79(3341)

18 cans of soda for $9
9.00/18=$.50 PER CAN BEST BUY IS METROMARKET.
10 cans of soda for $7.50
7.50/10=$.75 PER CAN
14 cans of soda for $7.35
7.35/14=$.525 PER CAN.